Calculate the magnitude of the linear momentum for the following cases.


(a) a proton with mass 1.67  10-27 kg, moving with a speed of 5.85  106 m/s

 ___kg · m/s


(b) a 14.5-g bullet moving with a speed of 455 m/s

 ___kg · m/s


(c) a 74.0-kg sprinter running with a speed of 12.5 m/s

 ___kg · m/s


(d) the Earth (mass = 5.98  1024 kg) moving with an orbital speed equal to 2.98  104 m/s.

___ kg · m/s



1
Expert's answer
2021-09-30T18:27:49-0400

(a) a proton with mass 1.67  10-27 kg, moving with a speed of 5.85  106 m/s:


"p=mv=(1.67\times 10^{-27}kg)( 5.85\times 10^6 \frac{m}{s})"

p = 9.7695 X 10-21 kg · m/s


(b) a 14.5-g bullet moving with a speed of 455 m/s:


"p=mv=(14.5\times 10^{-3}kg)( 455 \frac{m}{s})"

p = 6.5975 kg · m/s


(c) a 74.0-kg sprinter running with a speed of 12.5 m/s:


"p=mv=(74\,kg)( 12.5\, \frac{m}{s})"

p = 925 kg · m/s


(d) the Earth (mass = 5.98  1024 kg) moving with an orbital speed equal to 2.98  104 m/s:


"p=mv=(5.98\times 10^{24}kg)( 2.98\times 10^4 \frac{m}{s})"

p = 1.78204 X 1029 kg · m/s



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.